# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ V-B Proof of Theorem 2
We remark that if a standard matrix Bernstein inequality [80] were used in the proof of Lemma 7, the constants $c_{\alpha}$ and $c_{\gamma}$ would scale as $(\log n)^{1/2}$. From Eq. 63, we would get an additional factor of $(\log n)^{1/2}$ in the scaling of $\varepsilon^{2}_{\mathsf{c}}$, thus proving Proposition 4.
@@ V-C1 Proof of Lemma 7
$T_{1b}\triangleq\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}-2\left(\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)-\pi_{j}\left(\pi_{i}+\pi_{ij}\right)\right)\left(\hat{p}_{ij}-p_{ij}\right).$
@@
$\forall t_{1},t_{2}>0,\ {\mathbb{P}}(T_{1a}+T_{1b}\geq t_{1}+t_{2})\leq{\mathbb{P}}(T_{1a}\geq t_{1})+{\mathbb{P}}(T_{1b}\geq t_{2}).$
Hence, we proceed by bounding tail bounds on $T_{1a}$ and $T_{1b}$ separately as below. These bounds will be derived by conditioning that the event $\mathcal{A}_{2}$ holds, where we define
$\mathcal{A}_{2}\triangleq\bigg{\{}\|{\hat{\pi}}-\pi\|_{2}\leq c_{2}\frac{\|\pi\|_{2}}{\sqrt{kd_{\max}}}\bigg{\}}$ (64)
and by Lemma 6 we know that ${\mathbb{P}}(\mathcal{A}_{2})\geq 1-O(n^{-3})$.
Tail Bounds for $T_{1a}$: Define a matrix ${Q}$ as
@@
$=(\hat{\pi}-\pi)^{\mathrm{T}}{Q}(\hat{\pi}+\pi)+(\hat{\pi}-\pi)^{\mathrm{T}}{Q}^{\mathrm{T}}(\hat{\pi}+\pi)+\underbrace{\left({\hat{\pi}}^{2}-\pi^{2}\right)^{\mathrm{T}}{Q}\mathbf{1}_{n}+\mathbf{1}_{n}^{\mathrm{T}}{Q}\left(\hat{\pi}^{2}-\pi^{2}\right)}_{\zeta_{0}}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}2c_{2}\frac{\|\pi\|_{2}}{\sqrt{kd_{\max}}}\|Q\|_{2}(2\|\pi\|_{2}+\|{\hat{\pi}}-\pi\|_{2})+2\tilde{c}_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}4c_{2}n\|\pi\|_{\infty}^{2}\frac{\|Q\|_{2}}{\sqrt{kd_{\max}}}+2c_{2}^{2}n\|\pi\|^{2}_{\infty}\frac{\|Q\|_{2}}{kd_{\max}}+2\tilde{c}_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}$
$\leq\hat{c}_{2}n\|\pi\|_{\infty}^{2}\frac{\|Q\|_{2}}{\sqrt{kd_{\max}}},$ (65)
where $\zeta_{1}$ follows from the fact $x^{\mathrm{T}}Ay\leq\|x\|_{2}\|y\|_{2}\|A\|_{2}$ and since the event $\mathcal{A}_{2}$ holds. In $\zeta_{2}$, we utilize the fact that $\|\pi\|_{2}\leq\sqrt{n}\|\pi\|_{\infty}$. The term $\zeta_{0}$ is upper bounded as
$\zeta_{0}\leq 2\tilde{c}_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}.$ (66)
@@
$\left(\hat{\pi}^{2}-\pi^{2}\right)^{\mathrm{T}}Q{\mathbf{1}}_{n}$ $=(\hat{\pi}-\pi)^{\mathrm{T}}(\hat{\Pi}+\Pi)Q{\mathbf{1}}_{n}=(\hat{\pi}-\pi)^{\mathrm{T}}(\hat{\Pi}-\Pi+2\Pi)Q{\mathbf{1}}_{n}$
@@
$\leq\frac{c_{2}^{2}\|\pi\|_{2}^{2}}{kd_{{\max}}}\left\|Q{\mathbf{1}}_{n}\right\|_{\infty}+\frac{2c_{2}\|\pi\|_{2}}{\sqrt{kd_{{\max}}}}\|\pi\|_{\infty}\|Q\|_{2}\sqrt{n}$
@@
Now we will establish concentration bounds for $\|Q{\mathbf{1}}_{n}\|_{\infty}$. To accomplish this, we will utilize McDiamard’s inequality [80]. However, first note that
$\left(Q{\mathbf{1}}_{n}\right)_{i}=\sum_{j:(i,j)\in{\mathcal{E}}}\left(\hat{Y}_{ij}-p_{ij}^{2}\right)=\sum_{j:(i,j)\in{\mathcal{E}}}\frac{\left(\sum_{m=1}^{k}Z_{m,{ij}}\right)\left(\sum_{m=1}^{k}Z_{m,{ij}}-1\right)}{k(k-1)}-p_{ij}^{2}.$
@@
$V_{ij}=\frac{\left(\sum_{m=1}^{k}Z_{m,{ij}}\right)\left(\sum_{m=1}^{k}Z_{mij}-1\right)}{k(k-1)}.$
Let $V_{ij}^{\prime}$ be the value of $V_{ij}$ when one of $Z_{m_{ij}}$ is replaced by $Z_{m,{ij}}^{\prime}$, i.e., we have
@@
$\left|V^{\prime}_{ij}-V_{ij}\right|$ $=\left|\frac{2\left(Z_{m,{ij}}^{\prime}-Z_{m,{ij}}\right)Z_{ij}}{k(k-1)}+\frac{\left(Z_{m,{ij}}^{\prime}-Z_{m,{j}j}\right)\left(Z_{m,{ij}}^{\prime}-Z_{m,{ij}-1}\right)}{k(k-1)}\right|$
$\leq\frac{2|Z_{m,{ij}}^{\prime}-Z_{m_{ij}}|}{k(k-1)}\left|Z_{ij}+\frac{(Z_{m,{ij}}^{\prime}-Z_{m,{ij}}-1)}{2}\right|\leq\frac{2}{k-1}.$
@@
$\mathbb{P}\left(\bigg{|}\sum_{j:(i,j)\in{\mathcal{E}}}\hat{Y}_{ij}-p_{ij}^{2}\bigg{|}>t\right)\leq 2\exp\left(\frac{-2t^{2}}{kd_{\max}\left(\frac{2}{k-1}\right)^{2}}\right).$
Substituting $t=c\sqrt{\frac{d_{\max}\log n}{k}}$, for some constant $c$, we obtain the following bound
$\forall\ i\in[n],\ {\mathbb{P}}\bigg{(}\left(Q{\mathbf{1}}_{n}\right)_{i}\geq c\sqrt{\frac{d_{\max}\log n}{k}}\bigg{)}\leq O(n^{-4}).$
Therefore, using union bound we have
${\mathbb{P}}\bigg{(}\left\|Q{\mathbf{1}}_{n}\right\|_{\infty}\geq c\sqrt{\frac{d_{\max}\log n}{k}}\bigg{)}\leq\sum_{i=1}^{n}{\mathbb{P}}\bigg{(}\left(Q{\mathbf{1}}_{n}\right)_{i}\geq c\sqrt{\frac{d_{\max}\log n}{k}}\bigg{)}\leq O(n^{-3}).$ (68)
Combining Eq. 67 and Eq. 68, and utilizing the fact that $d_{\max}\geq\log n$
$\left(\pi^{2}-\pi^{2}\right)^{\mathrm{T}}Q{\mathbf{1}}_{n}\leq cc_{2}^{2}\frac{n\|\pi\|^{2}_{\infty}}{k^{3/2}}\sqrt{\frac{\log n}{d_{\max}}}+2c_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}\leq\tilde{c}_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}.$
@@
Bounding $\|Q\|_{2}$: Now it remains to show that $\|Q\|_{2}\leq O(\sqrt{d_{\max}/k})$ with high probability. In the case of a complete graph, it is much easier to show because each entry $Q_{ij}$ is bounded, and therefore ${Q}$ is a random sub-Gaussian matrix, and the variance of each entry is upper bounded by $4/k$ (by Eq. 52). Hence, by [80, Theorem 4.4.5], the spectral norm $\|{Q}\|_{2}\leq 2c_{q}(2\sqrt{d_{\max}}+t)/\sqrt{k}$ for some constant $c_{q}$ with probability at least $1-2e^{-t^{2}}$. Substituting $t=\sqrt{\log n}$, we get the following bound with a probability at least $1-O(1/n^{3})$
@@
For a general graph model (with $d_{\max}\geq\log n$), an application of matrix Bernstein inequality yields $\|Q\|_{2}\leq O(\sqrt{\frac{d_{\max}\log n}{k}})$ (with high probability). The extra $\log n$ factor becomes a bottleneck later in the analysis. However, using recent advances in concentration inequalities [81, Corollary 2.15] we can indeed show that if $d_{\max}\geq(\log n)^{4}$, then $\|Q\|_{2}\leq O(\sqrt{d_{\max}/k})$ with high probability. The tail bounds of $\|Q\|_{2}$ are computed in Lemma 15, and therefore, we have
$\mathbb{P}\left(\|Q\|_{2}\geq\sqrt{\frac{24d_{\max}}{k}}+c\left(\frac{d_{\max}^{1/4}}{\sqrt{k}}(\log n)^{3/4}+\sqrt{\frac{t}{k}}+\frac{d_{\max}^{1/3}t^{2/3}}{(k^{2}(k-1))^{1/3}}+\frac{t}{k(k-1)}\right)\right)\leq 4ne^{-t}.$ (70)
Tail Bounds for $T_{1b}$: Now we bound $T_{1b}$ as
@@
$=\sum_{j=1}^{n}-2\left(\hat{\pi}_{j}^{2}-\pi_{j}^{2}\right)\left(\sum_{i:(i,j)\in{\mathcal{E}},i\neq j}(\hat{p}_{ij}-p_{ij})\right)-2\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\hat{\pi}_{i}-\pi_{i}\right)\left(\hat{\pi}_{j}-\pi_{j}\right)\left(\hat{p}_{ij}-p_{ij}\right)$
$\ \ \ \ \ \ \ \ \ \ -2\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\pi_{i}\left(\hat{p}_{ij}-p_{ij}\right)\left(\hat{\pi}_{j}-\pi_{j}\right)+\pi_{j}\left(\hat{p}_{ij}-p_{ij}\right)\left(\hat{\pi}_{i}-\pi_{i}\right)$
@@
$\leq\frac{4c_{2}n\left\|{\pi}\right\|_{\infty}^{2}}{\sqrt{kd_{\max}}}\|\hat{P}-P\|_{2}+\tilde{c}_{2}\frac{n\|\pi\|_{\infty}^{2}}{k^{1.5}}+2\|{\hat{\pi}}-\pi\|^{2}_{2}\|\hat{P}-P\|_{2}+4\|\pi\|_{2}\|\hat{P}-P\|_{2}\|\hat{\pi}-\pi\|_{2}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}\frac{4c_{2}n\left\|{\pi}\right\|_{\infty}^{2}}{\sqrt{kd_{\max}}}\|\hat{P}-P\|_{2}+\tilde{c}_{2}\frac{n\|\pi\|_{\infty}^{2}}{k^{1.5}}+2c_{2}^{2}n\|\pi\|^{2}_{\infty}\frac{\|\hat{P}-P\|_{2}}{kd_{\max}}+4c_{2}n\|\pi\|_{\infty}^{2}\frac{\|\hat{P}-P\|_{2}}{\sqrt{kd_{\max}}}$
$\leq\hat{c}_{2}n\|\pi\|_{\infty}^{2}\frac{\|\hat{P}-P\|_{2}}{\sqrt{kd_{\max}}}+\tilde{c}_{2}\frac{n\|\pi\|_{\infty}^{2}}{k^{1.5}},$ (71)
where $\zeta_{1}$ follows since the event $\mathcal{A}_{2}$ holds and the term $\zeta_{0}$ is bounded as follows and the justification is provided below
$|2\mathbf{1}_{n}^{\mathrm{T}}(\hat{P}-P)\left(\hat{\pi}^{2}-\pi^{2}\right)|\leq\frac{4c_{2}n\left\|{\pi}\right\|_{\infty}^{2}}{\sqrt{kd_{\max}}}\|\hat{P}-P\|_{2}+\tilde{c}_{2}\frac{n\|\pi\|_{\infty}^{2}}{k^{1.5}}.$ (72)
Bounding $\zeta_{0}$: To bound ${\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)\left(\hat{\pi}^{2}-\pi^{2}\right)$, we utilize the same trick as used in Eq. 66. Observe that
${\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)\left(\hat{\pi}^{2}-\pi^{2}\right)$ $={\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)(\hat{\Pi}+\Pi)(\hat{\pi}-\pi)={\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)(\hat{\Pi}-\Pi+2\Pi)(\hat{\pi}-\pi)$
@@
$\leq\|{\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)\|_{\infty}\|\hat{\pi}-\pi\|_{2}^{2}+2\sqrt{n}\|\hat{P}_{n}-P\left\|{}_{2}\right\|\pi\left\|{}_{\infty}\right\|\hat{\pi}-\pi\|_{2}$
$\leq\|{\mathbf{1}}_{n}(\hat{P}-P)\|_{\infty}\frac{c_{2}^{2}\|\pi\|_{2}^{2}}{kd_{\max}}+2\sqrt{n}\|\pi\|_{\infty}\|\hat{P}_{n}-P\|_{2}\frac{c_{2}\|\pi\|_{2}}{\sqrt{kd_{\max}}}$
$\leq\|{\mathbf{1}}_{n}(\hat{P}-P)\|_{\infty}\frac{c_{2}^{2}n\|\pi\|_{\infty}^{2}}{kd_{\max}}+2c_{2}\frac{n\|\pi\|_{\infty}^{2}}{\sqrt{kd_{\max}}}\|\hat{P}_{n}-P\|_{2}.$ (73)
@@
$({\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P))_{i}=\sum_{j:(i,j)\in{\mathcal{E}}}\hat{p}_{ji}-p_{ji}=\frac{1}{k}\sum_{j:(i,j)\in{\mathcal{E}}}\sum_{m=1}^{k}\left(Z_{m,{ji}}-p_{ji}\right).$
Utilizing Hoeffding’s inequality [80], we get
$P(|({\mathbf{1}}_{n}(\hat{P}-P))_{i}|>t)\leq 2\exp\bigg(\frac{-2t^{2}}{kd_{\max}(\frac{1}{k^{2}})}\bigg{missing}).$

### v2

@@ V-B Proof of Theorem 2
We remark that if a standard matrix Bernstein inequality [HDP] were used in the proof of Lemma 7, the constants $c_{\alpha}$ and $c_{\gamma}$ would scale as $(\log n)^{1/2}$. From Eq. 60, we would get an additional factor of $(\log n)^{1/2}$ in the scaling of $\varepsilon^{2}_{\mathsf{c}}$, thus proving Proposition 4.
@@ V-C1 Proof of Lemma 7
$T_{1b}\triangleq\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}-2\left(\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)-\pi_{j}\left(\pi_{i}+\pi_{j}\right)\right)\left(\hat{p}_{ij}-p_{ij}\right).$
@@
$\forall t_{1},t_{2}>0,\ {\mathbb{P}}(T_{1a}+T_{1b}$ $\geq t_{1}+t_{2})$
$\leq{\mathbb{P}}(T_{1a}\geq t_{1})+{\mathbb{P}}(T_{1b}\geq t_{2}).$
Hence, we proceed by bounding tail bounds on $T_{1a}$ and $T_{1b}$ separately as below. These bounds will be derived when the event $\mathcal{A}_{2}$ holds, where we define
$\mathcal{A}_{2}\triangleq\bigg\{\|{\hat{\pi}}-\pi\|_{2}\leq c_{2}\frac{\sqrt{n}\|\pi\|_{\infty}}{\sqrt{kd_{\max}}}\bigg\},$ (61)
where we have absorbed the parameters $h,\xi$ in a special constant $c_{2}$ for conciseness and we have $c_{2}=c\sqrt{h}/\xi^{2}$. Moreover, by Lemma 6, we know that ${\mathbb{P}}(\mathcal{A}_{2})\geq 1-O(n^{-3})$.
Tail Bounds for $T_{1a}$: Define a matrix Q{Q} as
@@
$=(\hat{\pi}-\pi)^{\mathrm{T}}{Q}(\hat{\pi}+\pi)+(\hat{\pi}-\pi)^{\mathrm{T}}{Q}^{\mathrm{T}}(\hat{\pi}+\pi)$
$\ \ \ \ \ \ \ +\underbrace{\left({\hat{\pi}}^{2}-\pi^{2}\right)^{\mathrm{T}}{Q}\mathbf{1}_{n}+\mathbf{1}_{n}^{\mathrm{T}}{Q}\left(\hat{\pi}^{2}-\pi^{2}\right)}_{\zeta_{0}}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}2c_{2}\frac{\sqrt{n}\|\pi\|_{\infty}}{\sqrt{kd_{\max}}}\|Q\|_{2}(2\|\pi\|_{2}+\|{\hat{\pi}}-\pi\|_{2})$
$\ \ \ \ \ \ \ +2{c}_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}+\tilde{c}{c}_{2}\frac{n\|\pi\|_{\infty}^{2}}{k}$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}4c_{2}n\|\pi\|_{\infty}^{2}\frac{\|Q\|_{2}}{\sqrt{kd_{\max}}}+2c_{2}^{2}n\|\pi\|^{2}_{\infty}\frac{\|Q\|_{2}}{kd_{\max}}$
$\ \ \ \ \ \ \ \ +2{c}_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}+\tilde{c}{c}_{2}\frac{n\|\pi\|_{\infty}^{2}}{k}$
$\leq 6{c}_{2}n\|\pi\|_{\infty}^{2}\frac{\|Q\|_{2}}{\sqrt{kd_{\max}}}+\tilde{c}{c}_{2}\frac{n\|\pi\|_{\infty}^{2}}{k},$ (62)
where $\zeta_{1}$ follows from the fact $x^{\mathrm{T}}Ay\leq\|x\|_{2}\|y\|_{2}\|A\|_{2}$ and since the event $\mathcal{A}_{2}$ holds. In $\zeta_{2}$, we utilize the fact that $\|\pi\|_{2}\leq\sqrt{n}\|\pi\|_{\infty}$, and the last inequality follows since $k\geq O(h/({d}\xi^{4}))$. The term $\zeta_{0}$ is upper bounded as
$\zeta_{0}\leq 2{c}_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}+\tilde{c}{c}_{2}\frac{n\|\pi\|_{\infty}^{2}}{k}.$ (63)
@@
$\left(\hat{\pi}^{2}-\pi^{2}\right)^{\mathrm{T}}Q{\mathbf{1}}_{n}$
$=(\hat{\pi}-\pi)^{\mathrm{T}}(\hat{\Pi}+\Pi)Q{\mathbf{1}}_{n}=(\hat{\pi}-\pi)^{\mathrm{T}}(\hat{\Pi}-\Pi+2\Pi)Q{\mathbf{1}}_{n}$
@@
$\leq\frac{c_{2}^{2}n\|\pi\|_{\infty}^{2}}{kd_{{\max}}}\left\|Q{\mathbf{1}}_{n}\right\|_{\infty}+\frac{2c_{2}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{kd_{{\max}}}}\|\pi\|_{\infty}\|Q\|_{2}\sqrt{n}$
@@
Now we will establish concentration bounds for $\|Q{\mathbf{1}}_{n}\|_{\infty}$. To accomplish this, we will utilize McDiarmid’s inequality [HDP]. However, first note that
$(Q$ ${\mathbf{1}}_{n})_{i}=\sum_{j:(i,j)\in{\mathcal{E}}}\left(\hat{Y}_{ij}-p_{ij}^{2}\right)$
$=\sum_{j:(i,j)\in{\mathcal{E}}}\frac{\left(\sum_{m=1}^{k}Z_{m,{ij}}\right)\left(\sum_{m=1}^{k}Z_{m,{ij}}-1\right)}{k(k-1)}-p_{ij}^{2}.$
@@
$V_{ij}=\frac{\left(\sum_{m=1}^{k}Z_{m,{ij}}\right)\left(\sum_{m=1}^{k}Z_{m,ij}-1\right)}{k(k-1)}.$
Let $V_{ij}^{\prime}$ be the value of $V_{ij}$ when one of $Z_{m,{ij}}$ is replaced by $Z_{m,{ij}}^{\prime}$, i.e., we have
@@
$\left|V^{\prime}_{ij}-V_{ij}\right|$ $=\bigg|\frac{2\left(Z_{m,{ij}}^{\prime}-Z_{m,{ij}}\right)Z_{ij}}{k(k-1)}$
$\ \ \ \ \ \ \ \ \ \ +\frac{\left(Z_{m,{ij}}^{\prime}-Z_{m,ij}\right)\left(Z_{m,{ij}}^{\prime}-Z_{m,{ij}-1}\right)}{k(k-1)}\bigg|$
$\leq\frac{2|Z_{m,{ij}}^{\prime}-Z_{m,{ij}}|}{k(k-1)}\left|Z_{ij}+\frac{(Z_{m,{ij}}^{\prime}-Z_{m,{ij}}-1)}{2}\right|$
$\leq\frac{2}{k-1}.$
@@
$\mathbb{P}\left(\left|\sum_{j:(i,j)\in{\mathcal{E}}}\hat{Y}_{ij}-p_{ij}^{2}\right|>t\right)\leq 2\exp\left(\frac{-2t^{2}}{kd_{\max}\left(\frac{2}{k-1}\right)^{2}}\right).$
Substituting $t=\tilde{c}\sqrt{\frac{d_{\max}\log n}{k}}$, for some constant $\tilde{c}$, we obtain the following bound:
$\forall\ i\in[n],\ {\mathbb{P}}\bigg(\left(Q{\mathbf{1}}_{n}\right)_{i}\geq\tilde{c}\sqrt{\frac{d_{\max}\log n}{k}}\bigg)\leq O(n^{-4}).$
Therefore, using the union bound, we have
${\mathbb{P}}\bigg(\left\|Q{\mathbf{1}}_{n}\right\|_{\infty}$ $\geq\tilde{c}\sqrt{\frac{d_{\max}\log n}{k}}\bigg)$
$\leq\sum_{i=1}^{n}{\mathbb{P}}\bigg(\left(Q{\mathbf{1}}_{n}\right)_{i}\geq\tilde{c}\sqrt{\frac{d_{\max}\log n}{k}}\bigg)$
$\leq O(n^{-3}).$ (65)
Combining Eq. 64 and Eq. 65, and utilizing the fact that $c_{2}=c\frac{\sqrt{h}}{\xi^{2}}$ and $\xi^{4}kd_{\max}\geq O(h\log n)$, we have the following bound with high probability:
$\left(\pi^{2}\!-\!\pi^{2}\right)^{\mathrm{T}}Q{\mathbf{1}}_{n}\!$ $\leq\tilde{c}c_{2}\frac{n\|\pi\|^{2}_{\infty}}{k}\sqrt{\frac{ch\log n}{\xi^{4}kd_{\max}}}+2c_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}$
$\leq\tilde{c}c_{2}\frac{n\|\pi\|^{2}_{\infty}}{k}+2{c}_{2}\frac{n\|\pi\|_{\infty}^{2}\|Q\|_{2}}{\sqrt{kd_{\max}}}.$ (66)
@@
Bounding $\|Q\|_{2}$: Now it remains to show that $\|Q\|_{2}\leq O(\sqrt{d_{\max}/k})$ with high probability. In the case of a complete graph, it is much easier to show because each entry $Q_{ij}$ is bounded, and therefore Q{Q} is a random sub-Gaussian matrix, and the variance of each entry is upper bounded by $4/k$ (by Eq. 50). Hence, by [HDP, Theorem 4.4.5], the spectral norm $\|{Q}\|_{2}\leq 2c_{q}(2\sqrt{d_{\max}}+t)/\sqrt{k}$ for some constant $c_{q}$ with probability at least $1-2e^{-t^{2}}$. Substituting $t=\sqrt{\log n}$, we get the following bound with a probability at least $1-O(1/n^{3})$
@@
For a general graph model (with $d_{\max}\geq\log n$), an application of matrix Bernstein inequality [HDP] yields $\|Q\|_{2}\leq O(\sqrt{{d_{\max}\log n}/{k}})$ (with high probability). The extra $\log n$ factor becomes a bottleneck later in the analysis. However, using recent advances in concentration inequalities [brailovskaya2022, Corollary 2.15] we can indeed show that if $d_{\max}\geq(\log n)^{4}$, then $\|Q\|_{2}\leq O(\sqrt{d_{\max}/k})$ with high probability. The tail bounds of $\|Q\|_{2}$ are computed in Lemma 15, and therefore, we have
$\mathbb{P}\bigg(\|Q\|_{2}$ $\geq\sqrt{\frac{24d_{\max}}{k}}+c\bigg(\frac{d_{\max}^{1/4}}{\sqrt{k}}(\log n)^{3/4}+\sqrt{\frac{t}{k}}$
$\ \ \ +\frac{d_{\max}^{1/3}t^{2/3}}{(k^{2}(k-1))^{1/3}}+\frac{t}{k(k-1)}\bigg)\bigg)\leq 4ne^{-t}.$ (68)
Substituting $t=c\log n$, we obtain the bound $\|Q\|_{2}\leq O(\sqrt{d_{\max}/k})$, which by Eq. 62 implies the bound on $T_{1a}$ as $T_{1a}\leq O(c_{2}n\|\pi\|_{\infty}^{2}/k)$.
Tail Bounds for $T_{1b}$: Now we bound the quantity $T_{1b}$ as
@@
$=\sum_{j=1}^{n}-2\left(\hat{\pi}_{j}^{2}-\pi_{j}^{2}\right)\left(\sum_{\begin{subarray}{c}i:(i,j)\in{\mathcal{E}},\\ i\neq j\end{subarray}}(\hat{p}_{ij}-p_{ij})\right)$
$\ \ \ \ \ \ \ \ -2\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\hat{\pi}_{i}-\pi_{i}\right)\left(\hat{\pi}_{j}-\pi_{j}\right)\left(\hat{p}_{ij}-p_{ij}\right)$
$\ \ \ \ \ \ \ \ \ \ -2\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\pi_{i}\left(\hat{p}_{ij}-p_{ij}\right)\left(\hat{\pi}_{j}-\pi_{j}\right)$
$\ \ \ \ \ \ \ \ \ \ -2\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\pi_{j}\left(\hat{p}_{ij}-p_{ij}\right)\left(\hat{\pi}_{i}-\pi_{i}\right)$
@@
$\leq\frac{4c_{2}n\left\|{\pi}\right\|_{\infty}^{2}}{\sqrt{kd_{\max}}}\|\hat{P}-P\|_{2}+\tilde{c}c_{2}\frac{n\|\pi\|_{\infty}^{2}}{k}+$
$\ \ \ \ \ \ 2\|{\hat{\pi}}-\pi\|^{2}_{2}\|\hat{P}-P\|_{2}+4\|\pi\|_{2}\|\hat{P}-P\|_{2}\|\hat{\pi}-\pi\|_{2}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}\frac{4c_{2}n\left\|{\pi}\right\|_{\infty}^{2}}{\sqrt{kd_{\max}}}\|\hat{P}-P\|_{2}+\tilde{c}c_{2}\frac{n\|\pi\|_{\infty}^{2}}{k}$
$\ \ \ \ \ +2c_{2}^{2}n\|\pi\|^{2}_{\infty}\frac{\|\hat{P}-P\|_{2}}{kd_{\max}}+4c_{2}n\|\pi\|_{\infty}^{2}\frac{\|\hat{P}-P\|_{2}}{\sqrt{kd_{\max}}}$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}\hat{c}c_{2}n\|\pi\|_{\infty}^{2}\frac{\|\hat{P}-P\|_{2}}{\sqrt{kd_{\max}}}+\tilde{c}c_{2}\frac{n\|\pi\|_{\infty}^{2}}{k},$ (69)
where $\zeta_{1}$ holds on the event $\mathcal{A}_{2}$ and the term $\zeta_{0}$ is bounded as
$|2\mathbf{1}_{n}^{\mathrm{T}}(\hat{P}$ $-P)\left(\hat{\pi}^{2}-\pi^{2}\right)|$
$\ \ \ \ \ \leq\frac{4c_{2}n\left\|{\pi}\right\|_{\infty}^{2}}{\sqrt{kd_{\max}}}\|\hat{P}-P\|_{2}+\tilde{c}c_{2}\frac{n\|\pi\|_{\infty}^{2}}{k}.$ (70)
Moreover, $\zeta_{2}$} follows, for some constant $\hat{c}$, since $\sqrt{c_{2}^{2}/(kd_{\max})}=\sqrt{ch/(kd_{\max}\xi^{4})}\leq O(1)$ by our assumption in Theorem 2.
Bounding $\zeta_{0}$: [^5] The $\zeta_{0}$ notation is overloaded. Note that we bound a different $\zeta_{0}$ term this time to the one we addressed earlier in this proof. To bound ${\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)\left(\hat{\pi}^{2}-\pi^{2}\right)$, we utilize the same technique as used in Eq. 63. Observe that
${\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)\left(\hat{\pi}^{2}-\pi^{2}\right)$
$={\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)(\hat{\Pi}+\Pi)(\hat{\pi}-\pi)$
$={\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P)(\hat{\Pi}-\Pi+2\Pi)(\hat{\pi}-\pi)$
@@
$\leq\|{\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}\!-\!P)\|_{\infty}\|\hat{\pi}-\pi\|_{2}^{2}+2\sqrt{n}\|\hat{P}-P\left\|{}_{2}\right\|\pi\left\|{}_{\infty}\right\|\hat{\pi}\!-\!\pi\|_{2}$
$\leq\|{\mathbf{1}}_{n}(\hat{P}-P)\|_{\infty}\frac{c_{2}^{2}n\|\pi\|_{\infty}^{2}}{kd_{\max}}\!+\!2\sqrt{n}\|\hat{P}-P\|_{2}\frac{c_{2}\sqrt{n}\|\pi\|^{2}_{\infty}}{\sqrt{kd_{\max}}}$
$\leq\|{\mathbf{1}}_{n}(\hat{P}-P)\|_{\infty}\frac{c_{2}^{2}n\|\pi\|_{\infty}^{2}}{kd_{\max}}+2c_{2}\frac{n\|\pi\|_{\infty}^{2}}{\sqrt{kd_{\max}}}\|\hat{P}-P\|_{2}.$ (71)
@@
$({\mathbf{1}}_{n}^{\mathrm{T}}(\hat{P}-P))_{i}$ $=\sum_{j:(i,j)\in{\mathcal{E}}}(\hat{p}_{ji}-p_{ji})$
$=\frac{1}{k}\sum_{j:(i,j)\in{\mathcal{E}}}\sum_{m=1}^{k}\left(Z_{m,{ji}}-p_{ji}\right).$
Utilizing Hoeffding’s inequality [HDP], we get
[…]